# Probing the cosmic sterile-neutrino background with IceCube

### v1

@@ II Model description and methods
To estimate the rate of high-energy starting event (HESE) in IceCube, we use the Monte Carlo sample of simulated events [^1] https://github.com/icecube/HESE-7-year-data-release/ from Ref. Abbasi et al. (2021). The input parameters used to calculate the HESE event spectrum are tabulated in Table 1, and a description of these parameters can be found in Ref. Abbasi et al. (2021). Of these, we only consider the astrophysical neutrino spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) as variables and fix other parameters to their typical values. We assume that the new physics does not affect the fluxes of conventional and prompt atmospheric neutrinos, as well as atmospheric muons. A detailed calculation where all parameters are optimized is computationally very expensive and beyond the scope of this work. The event spectrum and the subsequent log-likelihood are obtained by reweighting the Monte Carlo events from astrophysical neutrinos. As we also scan over the mixing angles, the reweighting accounts for the flavor dependence of the interactions.
@@
Table 1: The set of physical and nuisance parameters that determine the HESE event rate are tabulated. We have shown the best-fit parameters for single power law (SPL) obtained in Ref. Abbasi et al. (2021) and the values/range of parameters used in this work. For a description of these variables, see TABLE VI.1 in Ref. Abbasi et al. (2021)
@@
We follow the analysis of Ref. Abbasi et al. (2021) closely and present our results in the frequentist statistical methodology. We report the new physics parameters that are most likely to explain the HESE data and compare it with the single power-law. The best-fit parameters and the errors are obtained using the profile likelihood,
~$\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})=\max_{\bm{\eta}}\mathcal{L}(\bm{\theta,\eta})\cdot\Pi(\bm{\theta,\eta})$ (4)
@@
where $\bm{\theta}$ is the set of parameters of interest and $\bm{\eta}$ is the set of nuisance parameters. The likelihood function used is the effective likelihood defined in Ref. Argüelles et al. (2019). The constraint $\Pi(\bm{\theta,\eta})$ is constant for parameters for which no external data is available. We have used the Python implementation of SLSQP algorithm to minimize the negative of the log of the function and thus obtain the best-fit point $\hat{\bm{\theta}}$. The model parameter test statistic is defined as,
$\text{TS}(\bm{\theta})=-2\,\log\,\left(\dfrac{\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})}{\tilde{\mathcal{L}}^{\texttt{profile}}(\hat{\bm{\theta}})}\right).$ (5)
The results are presented using Wilks’ theorem with appropriate degrees of freedom Wilks (1938). To validate our method, we first evaluate the 68 % C.L. and 90 % C.L. intervals for single-power law and compare them with the contours in Ref. Abbasi et al. (2021). We only scan over the astrophysical flux parameters and keep other parameters fixed to their best-fit values. We find that on expected lines, the best-fit point agrees very well, and the errors are underestimated.
@@
The HESE sample only includes events for which the interaction vertex is inside the restricted fiducial volume. To enhance the sensitivity to Glashow resonance (GR) events, a larger fiducial volume is required, which is achieved by looking at PeV energy partially contained events (PEPE). As the Monte Carlo event sample contains the GR, we use a simple rescaling for exposure to estimate the GR event rate. We find that our estimates agree with the Extended Data Table 1 in Ref. Aartsen et al. (2021a). The likelihood function, i.e., $\mathcal{L}(\bm{\theta,\eta})$ in Equation 4, is modified to include GR event for which we assume a simple Poisson likelihood.
@@ III Searching for dips in spectrum
Table 2: The best-fit values of the parameters of interest for single power law (SPL) obtained in Ref. Abbasi et al. (2021) and the result from our analysis is tabulated.
@@
Figure 2: Scan of the profile likelihood (see Equation 5) for the two new physics parameters: mass of the sterile neutrino ($m_{s}$) and mass of the vector boson mediator ($M_{X}$), is shown in orange. The darker (lighter) region denotes the 68% C.L. (90 % C.L.) interval. Note that for each point, the mixing angles are fixed to their best-fit values, and all other parameters are optimized. We also show two approximate curves that result in dips at 300 TeV and 800 TeV due to interaction with a cosmic sterile neutrino background. The region labeled as NSI (pink) indicated the parameter space for which dips occur due to interaction with the active neutrino background only and the dip due to sterile neutrino is outside the HESE analysis bound.
We first look at the space of the new physics parameters – sterile neutrino mass ($m_{s}$) and the vector mediator mass ($M_{X}$), and present our results in Figure 2. As expected [^2] The Figure 2 should be compared with Fig. 2 in Chauhan and Mohanty (2018), we find a large degenerate parameter space that results in dips in the event spectrum. We also find an island for $M_{X}=5-10$ MeV, which is largely independent of the sterile neutrino mass. In this region, the dip due to sterile neutrino is below the 60 TeV threshold of the HESE sample, and there is a dip between 300-800 TeV due to the heavier active neutrino. This is similar to the case of neutrino secret interaction ($\nu$SI), which is considered in Ref. Ng and Beacom (2014). We also highlight the $m_{s}=2.6$ eV that is hinted in recent results Serebrov et al. (2021); Abbasi et al. (2022b).
Figure 3: Scan of the profile likelihood for the two new physics parameters: astrophysical normalization and spectral index. The darker (lighter) regions show the 68 % C.L. (90 % C.L.) intervals. The region in blue is obtained using single power law, and the region in orange is obtained by including Glashow Resonance and dips due to sterile neutrino background. Note that other physical parameters are optimized for each scan point. As we do not optimize over all of the parameters (see text for details), the intervals are slightly smaller than the ones reported in Ref. Abbasi et al. (2021).
Figure 4: Left: We show the best-fit SPL flux from Ref. Abbasi et al. (2021) (blue) and the best-fit flux from our analysis (orange). The binned reconstructed flux from the HESE sample Abbasi et al. (2021) and PEPE sample is shown in black and red, respectively. Right: The deposited energy spectrum of HESE events (black) is shown with our event prediction using the best-fit parameters for SPL with dips (orange) and SPL-only (blue). The background from atmospheric neutrinos and muons is the same for both curves. The shaded grey region represents the 60 TeV threshold below which the backgrounds dominate the event rate.
@@
We now look at the space of astrophysical neutrino flux parameters – spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$). The scan of the profile likelihood is presented in Figure 3. We find that the inclusion of the dips leads to lower values of the spectral index. The inclusion of the Glashow resonance event also disfavors large values of the spectral index. This is one of the key results of the paper. At present, with only 7.5 years of data, there is a large overlap between the preferred regions for SPL and SPL with dips. However, the trend is clearly noticeable. In the future, with additional data from other neutrino telescopes, we anticipate that the overlap will be resolved. This was recently demonstrated in Ref. Fiorillo and Bustamante (2023), where the authors search for a bump-like feature in the spectrum. We also look at single parameter profile likelihood, and the results have been presented in Figure 1. It is clear that the inclusion of absorption features leads to a lowering of the spectral index and thus reduces the tension in the HESE and Northern Tracks samples. A future systematic study would also include the impact of absorption on the Northern Tracks sample.
@@
In order to highlight the difference between the SPL best-fit and the best-fit parameters for SPL with dips, we show the neutrino flux prediction overlaid with the reconstructed flux from the IceCube collaboration in Figure 4. This plot only indicates of the position of dips in the neutrino spectrum. In our analysis, the likelihood is computed using the data from the deposited energy event spectrum. We have shown the predicted event spectrum from the two sets of best-fit parameters in Figure 4. As mentioned earlier, we assume that the atmospheric neutrinos are unaffected by the new physics, and the background predictions for both fits are the same. In a more detailed analysis, one may consider varying the atmospheric neutrino parameters. As there is a large overlap in the 68% C.L. interval of the two fits, we choose not to show the uncertainty bands in these figures.

### v2

@@ II.3 High Energy Starting Events (HESE)
To estimate spectrum of HESE in IceCube, we use the Monte Carlo sample of simulated events [^1] https://github.com/icecube/HESE-7-year-data-release/ from Ref. Abbasi et al. (2021a). The input parameters used to calculate the HESE event spectrum are tabulated in Table 1, and a description of these parameters can be found in Ref. Abbasi et al. (2021a). Of these, we only consider the astrophysical neutrino spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$) as variables and fix other parameters to their typical values. We assume that the new physics does not affect the fluxes of conventional and prompt atmospheric neutrinos, as well as atmospheric muons. A detailed calculation where all parameters are optimized is computationally very expensive and beyond the scope of this work. The event spectrum and the subsequent log-likelihood are obtained by reweighting the Monte Carlo events from astrophysical neutrinos. As we also scan over the mixing angles, the reweighting accounts for the flavor dependence of the interactions.
The IceCube collaboration uses the effective likelihood defined in Ref. Argüelles et al. (2019) for estimating the physics parameters and their confidence and credible intervals Abbasi et al. (2021a). We follow their analysis closely, and present our results in the frequentist statistical methodology. For completeness, the likelihood analysis is summarised here and we refer the reader to Ref. Abbasi et al. (2021a) and Ref. Argüelles et al. (2019) for additional details. To handle the large MC statistical uncertainty due to small sample size in some bins, Ref. Argüelles et al. (2019) introduces a modified Poisson likelihood function (called effective likelihood),
$\mathcal{L}_{\text{eff}}(\mu,\sigma,k)=\left(\dfrac{\mu}{\sigma^{2}}\right)^{\frac{\mu^{2}}{\sigma^{2}}+1}\,\Gamma\left(k+\dfrac{\mu^{2}}{\sigma^{2}}+1\right)\left[k!\left(1+\dfrac{\mu}{\sigma^{2}}\right)^{k+\frac{\mu^{2}}{\sigma^{2}}+1}\Gamma\left(\dfrac{\mu^{2}}{\sigma^{2}}+1\right)\right]^{-1},$ (4)
where $k$ is the observed number of events, $\mu$ is the expected number of events, $\sigma$ is the MC statistical uncertainty for the bin. The likelihood function is given by,
$\mathcal{L}({\bm{\theta,\eta}})=\prod_{j}^{N_{\text{bins}}}\,\mathcal{L}_{\text{eff}}(\mu_{j}({\bm{\theta,\eta}}),\sigma_{j}({\bm{\theta,\eta}}),k_{j})$ (5)
where $N_{\text{bins}}$ is the number of bins, $\bm{\theta}$ is the set of parameters of interest, and $\bm{\eta}$ is the set of nuisance parameters. Note that, for a given analysis the set of nuisance parameters also includes physical parameters that are not being examined. In Ref. Abbasi et al. (2021a), the data and the MC events are first separated by their morphology (tracks, cascades, double-cascades) and then binned further in two observables. The tracks and cascades are binned in their reconstructed energy and reconstructed zenith angle, whereas the double cascades are binned in reconstructed energy and their reconstructed separation between cascades. The bin widths are chosen to be comparable to the detector resolution and there are 840 bins in observable quantities. These bins contain contributions from atmospheric as well as astrophysical events.
@@
Table 1: The set of physical and nuisance parameters that determine the HESE spectrum are tabulated. We have shown the best-fit parameters for single power law (SPL) obtained in Ref. Abbasi et al. (2021a) and the values/range of parameters used in this work. For a description of these variables, see TABLE VI.1 in Ref. Abbasi et al. (2021a)
@@
We follow the analysis of Ref. Abbasi et al. (2021a) closely to estimate the new physics parameters that are most likely to explain the HESE data, and compare it with the single power-law. The dips in the astrophysical neutrino spectrum will not affect the events from atmospheric neutrinos and muons. Moreover, the atmospheric contribution to the starting events in IceCube is primarily determined be the low-energy events and the zenith-angle distributions. Hence, in our analysis we will fix the parameters associated with atmospheric events to their nomianal values [^2] We have verified that keeping the atmospheric parameters free does not change the results significantly. This is primarily because the atmospheric parameters are determined by the lower energy events and their zenith-angle distributions., and only analyse astrophysical parameters (i.e., $\gamma_{\texttt{astro}}$ and $\Phi_{\texttt{astro}}$). Due to this approximation, the uncertainities are underestimated and a comparison in presented in Appendix A. The best-fit parameters and the confidence intervals are obtained using the profile likelihood,
$\tilde{\mathcal{L}}^{\texttt{profile}}(\bm{\theta})=\max_{\bm{\eta}}\mathcal{L}(\bm{\theta,\eta})\cdot\Pi(\bm{\theta,\eta})$ (6)
@@
where $\Pi(\bm{\theta,\eta})$ denote the constraints on parameters. For parameters for which no external data is available, $\Pi$ is constant. We have used the Python implementation of SLSQP algorithm to minimize the negative of the log of the function and thus obtain the best-fit point $\hat{\bm{\theta}}$. The model parameter test statistic is defined as,
[…]
The results are presented using Wilks’ theorem with appropriate degrees of freedom Wilks (1938).
II.4 Glashow Resonance
@@
The HESE sample only includes events for which the interaction vertex is inside the restricted fiducial volume. To enhance the sensitivity to Glashow resonance (GR) events, a larger fiducial volume is required, which is achieved by looking at PeV energy partially contained events (PEPE). The Monte Carlo simulation provided with the HESE data release can be utilised to predict the number of GR events in the PEPE sample by a simple rescaling. Due to the larger effective area, the expected event rate for the PEPE selection is about twice that of the HESE selection Aartsen et al. (2021a). However, the livetime of PEPE sample is only 4.6 years as opposed to HESE for which the livetime is 7.22 years. If $\lambda_{\text{HESE}}$ is the event rate in the GR energy bin (i.e., 5.6 PeV to 7.7 PeV ), obtained from the simulation, then one can estimate the expected number of events in the PEPE sample as,
$\lambda_{\text{PEPE}}(\bm{\theta,\eta})=\lambda_{\text{HESE}}(\bm{\theta,\eta})\times 2\times(4.6/7.22)$ (8)
We have verified that the estimates obtained using this method agree very well with the Extended Data Table 1 in Ref. Aartsen et al. (2021a). As only one event has been observed in this bin, we define the likelihood assuming Poisson statistics as,
$\mathcal{L}_{\text{GR}}(\bm{\theta,\eta})\approx\lambda_{\text{PEPE}}(\bm{\theta,\eta})\,e^{-\lambda_{\text{PEPE}}(\bm{\theta,\eta})}.$ (9)
The likelihood function, i.e., $\mathcal{L}(\bm{\theta,\eta})$ in Equation 5, is multiplied by $\mathcal{L}_{\text{GR}}(\bm{\theta,\eta})$ to obtain the combined fit. The impact of including the GR event in our analysis is shown in Appendix A.
II.5 Northern Tracks
The Northern Tracks analysis by IceCube focusses on measuring the flux of high-energy $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ from atmospheric and astrophysical sources. The event selection criterion samples muon-like [^3] We refer to both $\mu^{-}$ and $\mu^{+}$ as “muons” tracks originating from the Northern celestial hemisphere, which significantly reduces the background of atmospheric muons and results in high-purity neutrino-induced events. The sample contains mostly through-going muon tracks and a small contribution from starting tracks. Unlike the HESE sample, a detailed MC simulation and likelihood analysis for Northern Tracks is not publicly available. In order to estimate the impact of dips in the neutrino spectrum on Northern Tracks, we use a “theorist’s approach” Kistler and Beacom (2006); Laha et al. (2013); Gaisser et al. (2016) to calculate the event rates.
The measurable muon spectrum for Northern Tracks is given by,
$\dfrac{dN_{\text{NT}}}{dE_{\mu}^{\text{true}}}=\dfrac{dN_{\text{thru. tr.}}}{dE_{\mu}^{\text{true}}}+\dfrac{dN_{\text{st. tr.}}}{dE_{\mu}^{\text{true}}}$ (10)
where the first term is the contribution from the through-going tracks that originate outside the detector volume, and the second term denotes the contribution of the starting tracks for which the primary interaction vertex is inside the detector volume. We have used $E_{\mu}^{\text{true}}$ to denote the true energy of the muon which is different from the reconstructed muon energy. The muon spectrum from the through-going events is calculated using,
$\dfrac{dN_{\text{thru. tr.}}}{dE_{\mu}^{\text{true}}}=\sum_{\nu_{\mu},\bar{\nu}_{\mu}}\dfrac{T\,N_{A}}{\alpha+\beta E_{\mu}^{\text{true}}}\int d\Omega\int_{E_{\mu}^{\text{true}}}^{E_{\nu}^{\text{max}}(\theta_{z})}dE_{\nu}\,A(E_{\nu},\theta)e^{-\uptau(E_{\nu},\theta_{z})}\,\dfrac{d\Phi_{\nu}}{dE_{\nu}}\int_{0}^{1-E_{\mu}^{\text{true}}/E_{\nu}}dy\,\dfrac{d\sigma_{\nu}(E_{\nu},y)}{dy},$ (11)
[…]
The evaluation of Equation 11 is not only computationally expensive but also require input from IceCube about effective areas. We can simplify the expression by first assuming that $A(E_{\nu},\theta)\approx A=1\,\text{km}^{2}$. We also assume that $E_{\nu}^{\text{max}}(\theta_{z})\rightarrow\infty$. As a result, the expression can be simplified as,
$\dfrac{dN_{\text{thru. tr.}}}{dE_{\mu}^{\text{true}}}=\dfrac{2\pi\,T\,N_{A}\,A}{\alpha+\beta E_{\mu}^{\text{true}}}\int_{E_{\mu}^{\text{true}}}^{\infty}dE_{\nu}\,\mathcal{S}_{\text{NT}}(E_{\nu})\,\tilde{\sigma}(E_{\mu}^{\text{true}},E_{\nu})\,\dfrac{d\Phi_{\nu+\bar{\nu}}}{dE_{\nu}}$ (12)
where,
$\tilde{\sigma}(E_{\mu}^{\text{true}},E_{\nu})=\int_{0}^{1-E_{\mu}^{\text{true}}/E_{\nu}}dy\,\dfrac{1}{2}\,\left[\dfrac{d\sigma_{\nu}(E_{\nu},y)}{dy}+\dfrac{d\sigma_{\bar{\nu}}(E_{\nu},y)}{dy}\right]$ (13)
and the average attenuation (or shadow) factor for Northern Tracks is,
$\mathcal{S}_{\text{NT}}(E_{\nu})=\int_{-1}^{0}d(\cos\theta_{z})\,\dfrac{1}{2}\,\left[e^{-\uptau(E_{\nu},\theta_{z})}+e^{-\uptau(E_{\bar{\nu}},\theta_{z})}\right].$ (14)
Note that in Equation 12, the flux accounts for both $\nu_{\mu}$ and $\bar{\nu}_{\mu}$. We have neglected the contributions from leptonic interaction channels such as $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{\mu}+\mu^{-}$ as their cross-sections are significanly smaller than DIS.
For the starting-tracks contribution, we have followed Ref. Mena et al. (2014) and use the same approximations to obtain,
$\dfrac{dN_{\text{st. tr.}}}{dE_{\mu}^{\text{true}}}=2\pi\,T\,N_{A}\,\rho\,V\int_{0}^{\infty}dE_{\nu}\,\mathcal{S}_{\text{NT}}(E_{\nu})\,\dfrac{d\sigma(E_{\mu},E_{\nu})}{dE_{\mu}}\,\dfrac{d\Phi_{\nu+\bar{\nu}}}{dE_{\nu}}$ (15)
where the differential cross-section $d\sigma/dE_{\mu}$ is averaged over neutrino and anti-neutrino, $\rho=0.9167$ g/cm3 is the density of ice, and we take $V=1\,\text{km}^{3}$.
Lastly, we need to account for the detector resolution. The true energy deposited by the travelling muon ($E_{\mu}^{\text{true}}$) and the proxy muon energy that is reconstructed do not usually have a one-to-one mapping. We model this uncertainty as a Gaussian, and the measured muon spectrum is given by Mena et al. (2014),
$\dfrac{dN}{dE_{\mu}}=\int_{0}^{\infty}dE_{\mu}^{\text{true}}\,\dfrac{dN}{dE_{\mu}^{\text{true}}}\,\mathcal{G}\left(E_{\mu},E_{\mu}^{\text{true}},\sigma(E_{\mu}^{\text{true}})\right)$ (16)
where we have considered $\sigma(E_{\mu}^{\text{true}})=0.125\,E_{\mu}^{\text{true}}$ in our analysis, similar to Ref. Mena et al. (2014). We understand that the starting track events have a better energy resolution than the through-going tracks, however, we do not make this difference. We use Equation 12 and Equation 16 to estimate the contribution of astrophysical neutrinos to the Northern Tracks. For atmospheric contribution, we use the results in Ref. Abbasi et al. (2025).
The observed data for Northern tracks is provided in Ref.Abbasi et al. (2022b) , which we use for our likelihood analysis. The IceCube collaboration uses both energy distribution and zenith-angle distribution of the observed events to calculate the likelihood. The atmospheric neutrino contribution to the Northern Tracks dominates at lower energies and has a very strong and characteristic zenith-angle dependence. In this work, we only consider the muon-energy bins above 100 TeV to minimize the effect of atmospheric neutrinos. We do not consider binning in zenith-angle as the atrophysical neutrinos have almost flat distribution. We define the likelihood assuming Poisson distribution as,
$\mathcal{L}_{i}(\bm{\theta},\bm{\eta})=\frac{\lambda_{i}(\bm{\theta},\bm{\eta})^{k}e^{-\lambda_{i}(\bm{\theta},\bm{\eta})}}{k!}$ (17)
where $\lambda_{i}(\bm{\theta},\bm{\eta})$ is the expected number of events in the $i$-th bin obtained by integrating Equation 16, and $k$ is the number of number of observed events in the bin. In our analysis, we only consider the bins for which events are recorded and we do not consider the zenith-angle distribution of events. The likelihood, profile likelihood, and the test statistic are defined similar to Equation 5, Equation 6, and Equation 7, respectively.
II.6 Enhanced Starting-Track Event Selection (ESTES)
The Enhanced Starting-Track Event Selection (ESTES) in IceCube is a event selection criterion that focuses on identifying $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ interactions that originate inside the detector volume from all directions (northern sky and southern sky) Silva (2021); Silva and Mancina (2020); Mancina and Silva (2020, 2021); Abbasi et al. (2021b); Silva et al. (2023); Silva (2023); Abbasi et al. (2024). By employing a dynamical veto volume and neutrino self-veto, the backgrounds from atmospheric muons and atmospheric neutrinos can be efficiently rejected, especially from the southern sky. This allows IceCube to probe for astrophysical muon neutrinos down to 1 TeV. ESTES events benefit from good directional resolution and improved neutrino energy reconstruction compared to other event selections, and have enchanced sensitivity to southern-sky neutrino sources, including the Galactic plane. These event can also be used to study the diffuse flux of neutrinos. Assuming a single power-law, the best fit spectral index for astrophysical neutrinos is,
$\gamma^{\text{ESTES}}_{\texttt{astro}}=2.58^{+0.10}_{-0.09},$ (18)
which lies between the fits obtained from HESE and Northern Tracks.
In order to estimate the impact of dips on the ESTES data, we again adopt a theorist’s approach to event rates. The muon spectrum for starting-tracks can be obtained using Equation 15 and Equation 16. As the ESTES sample looks at events from all directions, the shadow factor can be approximated as,
$\mathcal{S}_{\text{ESTES}}(E_{\nu})=1+\mathcal{S}_{\text{NT}}(E_{\nu}),$ (19)
assuming that the events from the southern-sky are not attenuated. Our approximation for the energy resolution should work for ESTES data as wellAbbasi et al. (2024).
Similar to Northern Tracks, the IceCube collaboration uses both muon-energy and zenith-angle distribution for their statistical analysis. In this work, we will only only consider the muon-energy bins above 70 TeV for the likelihood analysis of the ESTES data. The bin likelihood, total likelihood, profile likelihood, and the test statistic are defined similar to Equation 17, Equation 5, Equation 6, and Equation 7, respectively.
@@ III.1 HESE and GR Analysis
Figure 2: Left: Scan of the profile likelihood (see Equation 7) for the two new physics parameters: mass of the sterile neutrino ($m_{s}$) and mass of the vector-boson mediator ($M_{X}$), is shown in orange. The darker (lighter) region denotes the 68% C.L. (90 % C.L.) interval. Note that for each point, the mixing angles are fixed to their best-fit values, and all other parameters are optimized. In this plot, we have fixed $g_{X}=0.1$. We also show two approximate curves that result in dips at 300 TeV and 800 TeV due to interaction with a cosmic sterile neutrino background. The region labeled as NSI (pink) indicated the parameter space for which dips occur due to interaction with the active neutrino background only and the dip due to sterile neutrino is outside the HESE analysis bound. Right: Same as the left figure but for $G_{X}=10^{6}\,G_{F}$
Table 2: The best-fit values of the parameters of interest for single power law (SPL) obtained in Ref. Abbasi et al. (2021a) and the result from our HESE+GR analysis is tabulated.
@@
[…]
@@
We now look at the space of astrophysical neutrino flux parameters – spectral index ($\gamma_{\texttt{astro}}$) and the normalization scale ($\Phi_{\texttt{astro}}$). The scan of the profile likelihood is presented in Figure 3. We find that the inclusion of the dips leads to lower values of the spectral index. The inclusion of the Glashow resonance event also disfavors large values of the spectral index. This is one of the key results of the paper. At present, with only 7.5 years of data, there is a large overlap between the preferred regions for SPL and SPL-with-dips. However, the trend is clearly noticeable. In the future, with additional data from other neutrino telescopes, we anticipate that the overlap will be resolved. This was recently demonstrated in Ref. Fiorillo and Bustamante (2023), where the authors search for a bump-like feature in the spectrum. We also look at single parameter profile likelihood, and the results have been presented in Figure 1. It is clear that due to the inclusion of dips in the spectrum, the fits to HESE and GR data prefer a harder spectral index.
@@
In order to highlight the difference between the SPL best-fit and the best-fit parameters for SPL-with-dips, we show the neutrino flux prediction overlaid with the reconstructed flux from the IceCube collaboration in Figure 4. This plot only indicates of the position of dips in the neutrino spectrum. In our analysis, the likelihood is computed using the data from the deposited energy event spectrum. We have shown the predicted event spectrum from the two sets of best-fit parameters in Figure 4. As mentioned earlier, we have verified that the atmospheric neutrino paramters are unaffected by the new physics, and the background predictions for both fits are statistically indistinguishable. In a more detailed analysis, one may consider varying the models for atmospheric neutrinos. As there is a large overlap in the 68% C.L. interval of the two fits, we choose not to show the uncertainty bands in these figures.
Figure 3: Scan of the profile likelihood for the two astrophysical neutrino flux parameters: normalization scale and spectral index obtained from analysis of HESE and GR data. The darker (lighter) regions show the 68 % C.L. (90 % C.L.) intervals. The region in blue is obtained using single power law, and the region in orange is obtained by including the dips due to sterile neutrino background. Note that other physical parameters are optimized for each point. As we do not optimize over all of the parameters (see text for details), the intervals are slightly smaller than the ones reported in Ref. Abbasi et al. (2021a). See Appendix A for comparison.
Figure 4: Left: We show the best-fit SPL flux from Ref. Abbasi et al. (2021a) (blue) and the best-fit flux from our analysis (orange). The binned reconstructed flux from the HESE sample Abbasi et al. (2021a) and PEPE Aartsen et al. (2021a) sample is shown in black and red, respectively. Right: The deposited energy spectrum of HESE events (black) is shown with our event prediction using the best-fit parameters for SPL with dips (orange) and SPL-only (blue). The background from atmospheric neutrinos and muons is the same for both curves. The shaded grey region represents the 60 TeV threshold below which the backgrounds dominate the event rate.
III.2 Northern Tracks Analysis
In Section II.5, we outlined our methodology for estimating the distribution of Northern Track events in IceCube. We now compare our predictions with observed IceCube data to extract the astrophysical flux normalization and spectral index. We find that our methods systematically underestimate the event rates. This is a significant limitation of our analysis, and our best-fit normalization differs substantially from the official IceCube results. A further limitation of our analysis is the treatment of the atmospheric backgrounds. The fit of atmospheric neutrino parameters are determined by the low-energy data as well as their zenith-angle distribution, and is largely independent of the dips. Subsequently, we adopt the atmospheric component as reported by IceCube, and only use the high-energy bins (>100>100\,TeV) for our likelihood analysis. Moreover, we do not utilise the zenith-angle distribution in our analysis. As a result of this simplified treatment of statistics, the uncertainities on our analysis are over-estimated. Therefore, our results should be interpreted in terms of trends and relative comparisons rather than as precise determinations of the flux parameters.
With these limitations in mind, we examine whether the Northern Track data can discriminate between two representative flux hypotheses, i.e., SPL and SPL-with-dips. For the SPL-with-dips analysis, we fix the New Physics parameters to the best-fit values from the HESE-only analysis (see Table 2). The results are presented in Figure 5, where we show the test-statistic for the two scenarios as well as our predicted event distributions. As can be seen, Northern Tracks are largely insensitive to the dips in the astrophysical neutrino spectrum. This insensitivity arises because the observed muons are predominantly through-going, and thus the observed muon energy does not correspond uniquely to the incident neutrino energy [^6] A neutrino with larger energy interacting further from the detector and a neutrino will lesser energy interacting closer to the detector lead to a through-going muon of similar energy.. As a result, any sharp spectral features are smeared out in the observed muon spectrum, and leave only a modest imprint on the event rates. The detector resolution (see Equation 16) leads to additional smearing for both hypotheses equally and is not the dominant cause of this insensitivity. As a result, the prediction for spectral index is indistinguishable for the two scenarios (see Figure 5, left). This is also evident in our comparison of the event distributions for the the best-fit values for the two hypothesis (see Figure 5, right). The event rates are statistically indistinguishable. The predicted normalizations for the two fits are slightly different. As mentioned earlier, our simplified treatment does not yield the correct normalisation and hence we do not report them. Due to the insensitivity of Northern Tracks data to the narrow spectral features, changing the location of the dips (i.e. a scan over $M_{X}$ and $m_{s}$) does not yield meaningful insights. On the other hand, broad spectral features (e.g., those considered in Fiorillo and Bustamante (2023) and Mohanty et al. (2019)) can lead to more pronounced differences and therefore are more testable with Northern Tracks.
Figure 5: Left: The test statistic (see Equation 7) for the single power-law fit of 9.5 years Northern Tracks sample Abbasi et al. (2022b) is shown in blue. In green, we show the test statistic after including absorption features using the best-fit New Physics parameters from Table 2. The horizontal grey dashed lines are used to denote the 68% C.L. and 90% C.L. intervals. Right: The event rates obtained using Equation 16 for the SPL and SPL+dips are shown in blue and green respectively. Note that the normalization for both fits is slightly different. The observed data points are taken from Abbasi et al. (2022b) and the errors are statistical only. The red shaded region is the atmospheric contribution. In the grey shaded region, the atmospheric contribution is large and hence excluded from our analysis (see text for details).
III.3 ESTES Analysis
In Section II.6, we outlined our methodology for estimating the distribution of ESTES events in IceCube. Similar to Northern Tracks analysis, we now compare our predictions with observed IceCube data to extract the astrophysical flux parameters as well as determine the sensitivity to dips in the spectrum. We again find that our simple methods systematically underestimate the ESTES event rates when compared to IceCube’s detailed simulations. Similar to northern tracks analysis, we fix the atmospheric neutrino background and use only the high energy bins (>70>70\,TeV) in the likelihood. Again, the results in this section should be interpreted as trends and relative comparisons rather than as precise determinations.
Our ESTES analysis is similar to Northern Tracks and the results are presented in Figure 6. Since these are track-like events that start inside the detector, there is a better relationship between the measured muon energy and the incident neutrino energy. As a result, any narrow spectral features lead to difference in the predicted spectral index (see Figure 6, left). The best fit values from our analysis are,
[…]
The lowering of the spectral index is comparable to the results from HESE-only analysis.
We also compare the event distributions for the the best-fit values for the two hypothesis in Figure 5 (right). The predicted normalizations for the two fits are different. In addition, we also show the event rates obtained by considering dips and the SPL best-fit values for spectral index and normalization for comparison.
Figure 6: Left: The test statistic (see Equation 7) for the single power-law fit of 10.5 years ESTES sample Abbasi et al. (2024) is shown in blue. In purple, we show the test statistic after including absorption features using the best-fit New Physics parameters from Table 2. The horizontal grey dashed lines are used to denote the 68% C.L. and 90% C.L. intervals. Right: The event rates obtained using Equation 16 for the SPL and SPL+dips are shown in blue and purple respectively. Note that the normalization for both fits is different. The purple dashed curve is obtained by considering dips with flux parameters from best-fit values of SPL, and shown only for comparison. The observed data points are taken from Abbasi et al. (2024) and the errors are statistical only. The red shaded region is the atmospheric contribution. In the grey shaded region, the atmospheric contribution is large and hence excluded from our analysis (see text for details).
III.4 Forecast for IceCube-Gen2
IceCube-Gen2 is the planned upgrade of the IceCube neutrino observatory, which is expected to have $\sim$10 times larger exposure to TeV–PeV neutrinos Aartsen et al. (2021b). To estimate the event rate in IceCube-Gen2, we scale the weights according to the increased effective area. The effective area for cascades in IceCube-Gen2 is given in Ref. Aartsen et al. (2021b). However, the effective area for tracks in IceCube-Gen2 is not publicly available at the time of writing this paper. In this work, we will assume that the effective area is the same for cascades and tracks. From the preceeding discussion, it is clear that starting events (HESE or ESTES) are good probes of dips in the spectrum. We have evaluated the HESE event rates using 25 years of IceCube-Gen2 exposure for the best-fit parameters from the SPL analysis and for the best-fit New Physics parameters from this work. The results are presented in Figure 7. The local significance of the dips can be tested with such large exposures. It must be noted that there are other proposed neutrino telescopes that will also be sensitive to the absorption features. A combined analysis will improve the discovery prospects Fiorillo and Bustamante (2023).